OpenStudy (anonymous):

-3x+6=-9 WILL GIVE MEDAL FOR CORRECT ANSWER!

OpenStudy (anonymous):

-3x+6=-9 -3x=-15 x=5

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

no problem